determine dimensions of the rectangle of a greatest area that can be inscribed
in a semi-circle of radius 15m

Answer by TimothyLamb(4379)

the max area rectangle is actually a square.
---
s = side of square
---
basic geometry leads to a right triangle with:
hypotenuse (h) = r = 15
adjacent side = s/2
opposite side = s
---
use pythagorean theorem:
---
h = sqrt( s^2 + (s/2)^2 )
h = sqrt( ss + (s/2)(s/2) )
h = sqrt( ss + (ss/4) )
h = sqrt( (4ss/4) + (ss/4) )
h = sqrt( (5ss/4) )
h = sqrt(5)(s/2)
s = 2h/sqrt(5)
s = 30/sqrt(5)
---
s = 13.416408 sq.m
---
max area = 180 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php